# Noisy Quantum Learning Theory

### v1

@@ D.1 Separating NISQ and NBQP
The fact that $\textsf{NISQ}\subseteq\textsf{NBQP}$ is immediate, since $\textsf{NBQP}_{\lambda}$ can already run any $\lambda$-noisy quantum circuit. Given a classical step in a NISQ algorithm, NBQP can perfectly simulate the classical computation on a $\lambda$-noisy quantum register for any constant $\lambda<\lambda_{\text{th}}$, where $\lambda_{\text{th}}$ is the threshold for a constant-rate fault-tolerant quantum error correction (FT-QEC) scheme as in [ABO97], with polynomial overhead. Now we will show that this inclusion is actually strict relative to an oracle, proving Theorem 2.3.
@@ D.1.1 Encoded Shuffling Simon’s Problem
The oracle we use to separate NISQ and NBQP is a modification of the Shuffling Simon’s Problem from [CCL23]. First we give some preliminary definitions.
@@
We will now use shuffling to construct a version of Simon’s problem that is challenging for a quantum device with bounded depth, as done in [CCL23], but for which an NBQP algorithm can learn the hidden secret even when noise is applied to the quantum shuffling oracle.
@@
Theorem 4.11 of [CCL23] gives a quantum algorithm that solves the Enc-$d$-SSP problem with a noiseless quantum circuit of depth $2d+1$. For any constant $\lambda<\lambda_{\textnormal{th}}$, a $\lambda$-noisy quantum circuit can use quantum fault-tolerance [ABO97] to simulate this noiseless algorithm with logarithmic overhead in the depth of the quantum circuit, recalling that the encoding scheme QEC can, with high probability, execute a circuit of depth $3d>2d+1$ without logical errors. ∎
In the remainder of this section, we show that no NISQ algorithm making polynomially many queries can solve Enc-d-SSP with probability $\geq 2/3$. The proof strategy will be to first show that any $\textsf{BPP}^{\textsf{QNC}_{d}}$ has a success probability exponentially small in $n$, following the argument in [CCL23]. Then we argue that for $d$ and a number of queries both polynomial in $n$, the outcome distribution of any NISQ algorithm can be simulated by a $\textsf{BPP}^{\textsf{QNC}_{d}}$ algorithm to vanishingly small total variation distance.
D.1.2 Bounding quantum success probability by classical combinatorics
We first prove necessary lemmas presented in [CCL23] for $d\textsf{-SSP}$, now for the encoded version of the problem. The definitions we present are largely adapted from [ACC+23, CCL23].
@@ D.1.2 Bounding quantum success probability by classical combinatorics
We defer the proof to Appendix G. This lemma is the crucial step that allows us to relate quantum success probability to classical combinatorics. The O2H bound tells us that the output distribution of a quantum algorithm making a bounded number of queries to the true oracle differs from one querying a shadow oracle which knows nothing about the crucial function $f$ by an amount controlled by the classical probability of picking a bitstring that lies in a smaller hidden wrapper nested within a larger one. Notably, this is completely independent of the FT-QEC scheme. Since every round of nesting makes the space bigger by a factor of $2^{n}$, O2H intuitively allows us to bound these difference between these output distributions by an exponentially small number. This intuition leads to the following bound on the success probability of an algorithm with limited quantum depth.
Theorem D.10 (Proven in Sections 5, 8 of [CCL23]).
@@
The formulation of $\textnormal{{Enc-}}d\textnormal{{-SSP}}$ assumes $\mathcal{F}\sim D(f,d)$. Moreover, in Appendix G we prove Lemma D.9, implying that the second condition of Theorem D.10 holds for $\mathcal{O}_{\mathcal{F}}=\mathcal{F}^{\textnormal{Enc}}$ defined in Eq. (26). Hence, we immediately obtain the following lemma.
@@ D.1.3 Reducing to NISQ
Start with $\mathcal{T}^{\prime}=\mathcal{T}$. Let $\bar{D}$ be a parameter representing a depth threshold in the noisy quantum circuit. If at node $u$ in $\mathcal{T}$, the noisy quantum circuit performs more than $\bar{D}$ queries to $O_{f,d}$, replace $u$ in $\mathcal{T}^{\prime}$ with a noiseless quantum circuit that measures the $n$-qubit maximally mixed state in the computational basis. If the circuit makes less than $\bar{D}$ queries, replace $u$ with the noiseless quantum circuit that simulates the $\lambda$-noisy circuit by simulating depolarizing noise at each layer. With $\mathcal{T}^{\prime}$ defined, we call upon the following lemma from [CCHL23]:
Lemma D.13 (In [CCHL23], Lemmas D.15 and D.16).
@@ D.2.1 Remark on criteria for noise-robust quantum learning algorithms
A more interesting challenge for a quantum learning algorithm arises when local physical errors on the oracle system, or during its coupling to the computer, can affect the encoded quantum information in a highly nonlocal way. If a few sparse errors on the query register only corrupt a bounded amount of the logical information we are trying to extract, then repetition and majority-vote type procedures can still be used to distill a high-fidelity estimate of the oracle’s behavior [BLM+23, IRY05, CRR05, HMdW03, SYNW06]. On the other hand, suppose an oracle loads a single computational-basis bitstring encoding a value between 0 and $2^{n}-1$. In this case, a single local error may be enough to completely corrupt the stored information, since an error on the $k$-th qubit can change the measured value by $2^{k-1}$. While this toy example, and the corresponding oracle we construct for our exponential separation, are deliberately designed to impede an NBQP learner, the underlying intuition is that noise-robust quantum learning theory often requires some notion of logical locality in how information is encoded.
@@ D.2.2 Hybrid argument
Lemma D.14 (Learning from similar oracles, Lemma B.4 in [CCHL23]).
@@
We find that the “lifted Simon’s oracle” used in [CCHL23] to separate NISQ from BQP suffices to separate NBQP as well.
@@
Lemma D.16 (Distinguishability of noisy output distributions, Lemma C.22 in [CCHL23]).
@@
Lemma D.17 (Distinguishability propagation in a learning tree, NBQP variant of Lemma B.2 in [CCHL23]).
@@ Appendix E Quantum Advantage in Noisy Purity Testing
A well-studied candidate for achieving quantum advantage in learning from entangled measurements is the problem of estimating the purity of a quantum state. Several works demonstrate that without at least $n$ additional qubits of quantum memory, the sample complexity of the simpler purity testing problem (distinguishing between a $n$-qubit Haar-random pure state and a maximally mixed state) scales exponentially with $n$ [ABO97, CCHL22, GHYZ24, CGY24]. On the other hand, running a simple SWAP test on two copies of the state a constant number of times suffices to distinguish these cases, because while the inner product of a pure state with itself is $1$, the output of a SWAP test on two copies of a maximally mixed state is exponentially close to zero. In Section E.1, we show that this supposed quantum advantage breaks down in the presence of order-1 local depolarizing noise.
Implicitly, this result demonstrates that the SWAP test, a crucial subroutine in many quantum-enhanced algorithms and learning/property-testing protocols [HBC+22, CSSC18, LMR14, BOW19, HM13, MdW18], is degraded in the presence of noise by a factor exponential in system size. Moreover, no adaptive strategy can be used to circumvent this fact. To recover quantum advantage when performing SWAP operations on entire copies of an uncharacterized quantum state or process, the underlying physical system must exhibit an intrinsic robustness to noise, effectively enabling a form of quantum error-correction. In Section E.2, we construct such a system using a tensor-network model for holographic duality known as the HaPPY code [PYHP15]. The purity testing problem is reformulated as detecting the presence of a bulk black hole microstate from noisy measurements of the dual boundary state.
@@ E.1 Breakdown of advantage in noisy two-copy purity testing
where $S_{x}$ is the sum over all permutation operators on $x$ copies of the $n$-qubit Hilbert space. From [CGY24, Lemma 16] we have
@@
$L(\ell)$ $\geq\left(1-\frac{4T^{2}}{2^{n}}\right)\prod_{t=1}^{T}\frac{\mathsf{tr}(F_{s}^{t}\cdot D_{\lambda}^{\otimes 2n}(S_{2}))}{\mathsf{tr}(F_{s}^{t})}\,.$ (37)
@@
$L(\ell)\geq\left(1-\frac{4T^{2}}{2^{n}}\right)\prod_{t=1}^{T}\frac{\mathsf{tr}(F_{s}^{t}\cdot D_{\lambda}^{\otimes 2n}(S_{2}))}{\mathsf{tr}(F_{s}^{t})}\geq\left(1-\frac{4T^{2}}{2^{n}}\right)\prod_{t=1}^{T}\frac{\mathsf{tr}(F_{s}^{t}\cdot D_{\lambda}^{\otimes 2n}[\rho_{S}])}{\mathsf{tr}(F_{s}^{t}\rho_{m})}\,$, (38)
@@
$\leq 2^{-2n+1}+2\sum_{s}\frac{\mathsf{tr}(F_{s}\mathcal{D}_{\lambda}^{\otimes 2n}[\textsf{SWAP}_{n}])^{2}}{\mathsf{tr}(F_{s})}\,.$ (42)
@@ E.2 Black hole detection in holographic duality
We instantiate this scenario using the HaPPY code [PYHP15], a holographic tensor-network code that implements an isometric bulk-boundary encoding and exhibits quantum error-correcting structure. In our construction, the inner bulk legs of the network encode the black hole degrees of freedom, all remaining bulk legs are fixed in a reference state, and the learner has noisy access only to boundary qubits. The task is then to distinguish, using only such noisy boundary measurements, whether the bulk black hole region is prepared in a single Haar-random pure state or in the maximally mixed state on the same Hilbert space. We now spell out this setup more precisely.
@@ E.2.1 Defining the task
The remainder of this section is dedicated to the proof of Theorem E.3. While we consider the erasure channel for a provable guarantee, we remark that the hexagonal HaPPY code exhibits numerical evidence of error thresholds against more general Pauli noise channels [HCM+20, FHMS21, JE21], and it is likely that the two-copy distinguishing procedure will remain efficient under these channels (including the local depolarizing noise model considered in the rest of this work), by using more sophisticated decoders than the one considered in our upper bound.
@@ E.2.2 Intractability with single copies
Consider the many-vs.-one distinguishing problem from Appendix E, i.e. distinguishing between an $n$-qubit state $\rho$ that is either maximally mixed or sampled from the Haar measure over pure states. [CCHL22] demonstrates that any noiseless algorithm without quantum memory (with a learning tree representation given by Definition F.3, without the noise channels), with access to individual copies of $\rho$, requires $\Omega(2^{n/2})$ samples to solve this discrimination problem with high probability. Note that the decision-version of our task of detecting the presence of a black hole can be reformulated precisely as the purity testing problem, with the distinction that the unknown state is Haar random over a logical subspace that is then mapped by isometries to a state on an exponentially larger physical Hilbert space. The crucial observation is that this task reduces to the information-theoretic lower bound obtained from a memoryless learning tree for standard purity testing.
@@ E.2.3 Noise-robust detection using joint measurements
As in the lower bound, we utilize a basic reduction to the standard purity-testing problem to establishes our $O(1)$ two-copy upper bound. The first step is to use the intrinsic error-correction properties of the tensor network. In particular, [PYHP15] proposes a simple hierarchical majority-vote strategy for decoding the boundary state, known as the greedy decoder. While less robust than other decoding strategies, the greedy decoder allows us to obtain rigorous decoding guarantees. In particular, [PYHP15] gives us the following lemma.
@@
Under an erasure channel of strength $\lambda$, each out-leg of the boundary state is corrupted independently with probability $\lambda$. [PYHP15] shows that when the greedy decoder is applied to $t$ radial layers of the tensor network, the probability that an erasure error has propagated to a particular leg at radius RtR-t is $\leq\frac{1}{12}(12\lambda)^{\varphi^{R-t}}$. Applying a union bound over all legs at radius $r$, which we have shown is
@@
Hence, given two copies of $\mathcal{E}_{\lambda}(\rho_{b})$, we can apply the greedy decoder to both, and with probability $\geq 1-O(\exp\exp(-r(c-1)))$, we obtain two copies of $\rho$. In our problem, $\rho$ is either maximally mixed or a Haar-random pure state $\lvert\psi\rangle\!\langle\psi\rvert$. Our problem is then reduced to the well-studied purity testing problem, for which there is a simple two-copy algorithm. One can simply run a SWAP test $\textsf{SWAP}(\rho,\rho)$ (e.g., as described in [MdW18]); on two arbitrary states $\rho,\sigma$, the test accepts with probability

### v2

@@ D.1 Separating NISQ and NBQP
The fact that $\textsf{NISQ}\subseteq\textsf{NBQP}$ is immediate, since $\textsf{NBQP}_{\lambda}$ can already run any $\lambda$-noisy quantum circuit. Given a classical step in a NISQ algorithm, NBQP can perfectly simulate the classical computation on a $\lambda$-noisy quantum register for any constant $\lambda<\lambda_{\text{th}}$, where $\lambda_{\text{th}}$ is the threshold for a constant-rate fault-tolerant quantum error correction (FT-QEC) scheme as in [aharonov1999faulttolerantquantumcomputationconstant], with polynomial overhead. Now we will show that this inclusion is actually strict relative to an oracle, proving Theorem 2.3.
@@ D.1.1 Encoded Shuffling Simon’s Problem
The oracle we use to separate NISQ and NBQP is a modification of the Shuffling Simon’s Problem from [Chia_2023]. First we give some preliminary definitions.
@@
We will now use shuffling to construct a version of Simon’s problem that is challenging for a quantum device with bounded depth, as done in [Chia_2023], but for which an NBQP algorithm can learn the hidden secret even when noise is applied to the quantum shuffling oracle.
@@
Theorem 4.11 of [Chia_2023] gives a quantum algorithm that solves the Enc-$d$-SSP problem with a noiseless quantum circuit of depth $2d+1$. For any constant $\lambda<\lambda_{\textnormal{th}}$, a $\lambda$-noisy quantum circuit can use quantum fault-tolerance [aharonov1999faulttolerantquantumcomputationconstant] to simulate this noiseless algorithm with logarithmic overhead in the depth of the quantum circuit, recalling that the encoding scheme QEC can, with high probability, execute a circuit of depth $3d>2d+1$ without logical errors. ∎
In the remainder of this section, we show that no NISQ algorithm making polynomially many queries can solve Enc-d-SSP with probability $\geq 2/3$. The proof strategy will be to first show that any $\textsf{BPP}^{\textsf{QNC}_{d}}$ has a success probability exponentially small in $n$, following the argument in [Chia_2023]. Then we argue that for $d$ and a number of queries both polynomial in $n$, the outcome distribution of any NISQ algorithm can be simulated by a $\textsf{BPP}^{\textsf{QNC}_{d}}$ algorithm to vanishingly small total variation distance.
D.1.2 Bounding quantum success probability with classical combinatorics
We first prove necessary lemmas presented in [Chia_2023] for $d\textsf{-SSP}$, now for the encoded version of the problem. The definitions we present are largely adapted from [arora2023quantum, Chia_2023].
@@ D.1.2 Bounding quantum success probability with classical combinatorics
We defer the proof to Appendix H. This lemma is the crucial step that allows us to relate quantum success probability to classical combinatorics. The O2H bound tells us that the output distribution of a quantum algorithm making a bounded number of queries to the true oracle differs from one querying a shadow oracle which knows nothing about the crucial function $f$ by an amount controlled by the classical probability of picking a bitstring that lies in a smaller hidden wrapper nested within a larger one. Notably, this is completely independent of the FT-QEC scheme. Since every round of nesting makes the space bigger by a factor of $2^{n}$, O2H intuitively allows us to bound these difference between these output distributions by an exponentially small number. This intuition leads to the following bound on the success probability of an algorithm with limited quantum depth.
Theorem D.10 (Proven in Sections 5, 8 of [Chia_2023]).
@@
The formulation of $\textnormal{{Enc-}}d\textnormal{{-SSP}}$ assumes $\mathcal{F}\sim D(f,d)$. Moreover, in Appendix H we prove Lemma D.9, implying that the second condition of Theorem D.10 holds for $\mathcal{O}_{\mathcal{F}}=\mathcal{F}^{\textnormal{Enc}}$ defined in Eq. (26). Hence, we immediately obtain the following lemma.
@@ D.1.3 Reducing to NISQ
Start with $\mathcal{T}^{\prime}=\mathcal{T}$. Let $\bar{D}$ be a parameter representing a depth threshold in the noisy quantum circuit. If at node $u$ in $\mathcal{T}$, the noisy quantum circuit performs more than $\bar{D}$ queries to $O_{f,d}$, replace $u$ in $\mathcal{T}^{\prime}$ with a noiseless quantum circuit that measures the $n$-qubit maximally mixed state in the computational basis. If the circuit makes less than $\bar{D}$ queries, replace $u$ with the noiseless quantum circuit that simulates the $\lambda$-noisy circuit by simulating depolarizing noise at each layer. With $\mathcal{T}^{\prime}$ defined, we call upon the following lemma from [nisq]:
Lemma D.13 (In [nisq], Lemmas D.15 and D.16).
@@ D.2.1 Remark on criteria for noise-robust quantum learning algorithms
A more interesting challenge for a quantum learning algorithm arises when local physical errors on the oracle system, or during its coupling to the computer, can affect the encoded quantum information in a highly nonlocal way. If a few sparse errors on the query register only corrupt a bounded amount of the logical information we are trying to extract, then repetition and majority-vote type procedures can still be used to distill a high-fidelity estimate of the oracle’s behavior [buhrman2022quantummajorityvote, iwama_2005, chakraborty_2005, H_yer_2003, suzuki2006robustquantumalgorithmsepsbiased]. On the other hand, suppose an oracle loads a single computational-basis bitstring encoding a value between 0 and $2^{n}-1$. In this case, a single local error may be enough to completely corrupt the stored information, since an error on the $k$-th qubit can change the measured value by $2^{k-1}$. While this toy example, and the corresponding oracle we construct for our exponential separation, are deliberately designed to impede an NBQP learner, the underlying intuition is that noise-robust quantum learning theory often requires some notion of logical locality in how information is encoded.
@@ D.2.2 Hybrid argument
Lemma D.14 (Learning from similar oracles, Lemma B.4 in [nisq]).
@@
We find that the “lifted Simon’s oracle” used in [nisq] to separate NISQ from BQP suffices to separate NBQP as well.
@@
Lemma D.16 (Distinguishability of noisy output distributions, Lemma C.22 in [nisq]).
@@
Lemma D.17 (Distinguishability propagation in a learning tree, NBQP variant of Lemma B.2 in [nisq]).
@@ Appendix E Quantum Advantage in Noisy Purity Testing
A well-studied candidate for achieving quantum advantage in learning from entangled measurements is the problem of estimating the purity of a quantum state. Several works demonstrate that without at least $n$ additional qubits of quantum memory, the sample complexity of the simpler purity testing problem (distinguishing between a $n$-qubit Haar-random pure state and a maximally mixed state) scales exponentially with $n$ [aharonov1999faulttolerantquantumcomputationconstant, chen2021exponentialseparationslearningquantum, gong2024samplecomplexitypurityinner, weiyuan_paulis]. On the other hand, running a simple SWAP test on two copies of the state a constant number of times suffices to distinguish these cases, because while the inner product of a pure state with itself is $1$, the output of a SWAP test on two copies of a maximally mixed state is exponentially close to zero. In Section E.1, we show that this supposed quantum advantage breaks down in the presence of order-1 local depolarizing noise.
Implicitly, this result demonstrates that the SWAP test, a crucial subroutine in many quantum-enhanced algorithms and learning/property-testing protocols [Huang_2022, Cincio_2018, Lloyd_2014, bădescu2017quantumstatecertification, Harrow:2012gwf, montanaro2018surveyquantumpropertytesting], is degraded in the presence of noise by a factor exponential in system size. Moreover, no adaptive strategy can be used to circumvent this fact. To recover quantum advantage when performing SWAP operations on entire copies of an uncharacterized quantum state or process, the underlying physical system must exhibit an intrinsic robustness to noise, effectively enabling a form of quantum error-correction. In Section E.2, we construct such a system using a tensor-network model for holographic duality known as the HaPPY code [happy_2015]. The purity testing problem is reformulated as detecting the presence of a bulk black hole microstate from noisy measurements of the dual boundary state.
@@ E.1 Breakdown of advantage in noisy two-copy purity testing
where $S_{x}$ is the sum over all permutation operators on $x$ copies of the $n$-qubit Hilbert space. From [weiyuan_paulis, Lemma 16] we have
@@
$L(\ell)$ $\geq\left(1-\frac{4T^{2}}{2^{n}}\right)\prod_{t=1}^{T}\frac{\mathsf{tr}(F_{s}^{t}\cdot\mathcal{D}_{\lambda}^{\otimes 2n}(S_{2}))}{\mathsf{tr}(F_{s}^{t})}\,.$ (37)
@@
$L(\ell)\geq\left(1-\frac{4T^{2}}{2^{n}}\right)\prod_{t=1}^{T}\frac{\mathsf{tr}(F_{s}^{t}\cdot\mathcal{D}_{\lambda}^{\otimes 2n}(S_{2}))}{\mathsf{tr}(F_{s}^{t})}\geq\left(1-\frac{4T^{2}}{2^{n}}\right)\prod_{t=1}^{T}\frac{\mathsf{tr}(F_{s}^{t}\cdot\mathcal{D}_{\lambda}^{\otimes 2n}[\rho_{S}])}{\mathsf{tr}(F_{s}^{t}\rho_{m})}\,$, (38)
@@
$\leq 2^{-2n+1}+2\sum_{s}\frac{\mathsf{tr}(F_{s}\mathcal{D}_{\lambda}^{\otimes 2n}[\textsf{SWAP}_{n}])^{2}}{2^{2n}\mathsf{tr}(F_{s})}\,.$ (42)
@@ E.2 Black hole detection in holographic duality
We instantiate this scenario using the HaPPY code [happy_2015], a holographic tensor-network code that implements an isometric bulk-boundary encoding and exhibits quantum error-correcting structure. In our construction, the inner bulk legs of the network encode the black hole degrees of freedom, all remaining bulk legs are fixed in a reference state, and the learner has noisy access only to boundary qubits. The task is then to distinguish, using only such noisy boundary measurements, whether the bulk black hole region is prepared in a single Haar-random pure state or in the maximally mixed state on the same Hilbert space. We now spell out this setup more precisely.
@@ E.2.1 Defining the task
The remainder of this section is dedicated to the proof of Theorem E.3. While we consider the erasure channel for a provable guarantee, we remark that the hexagonal HaPPY code exhibits numerical evidence of error thresholds against more general Pauli noise channels [numerics_happy_decoding_2020, Farrelly_2021, Jahn_2021], and it is likely that the two-copy distinguishing procedure will remain efficient under these channels (including the local depolarizing noise model considered in the rest of this work), by using more sophisticated decoders than the one considered in our upper bound.
@@ E.2.2 Intractability with single copies
Consider the many-vs.-one distinguishing problem from Appendix E, i.e. distinguishing between an $n$-qubit state $\rho$ that is either maximally mixed or sampled from the Haar measure over pure states. [chen2021exponentialseparationslearningquantum] demonstrates that any noiseless algorithm without quantum memory (with a learning tree representation given by Definition F.3, without the noise channels), with access to individual copies of $\rho$, requires $\Omega(2^{n/2})$ samples to solve this discrimination problem with high probability. Note that the decision-version of our task of detecting the presence of a black hole can be reformulated precisely as the purity testing problem, with the distinction that the unknown state is Haar random over a logical subspace that is then mapped by isometries to a state on an exponentially larger physical Hilbert space. The crucial observation is that this task reduces to the information-theoretic lower bound obtained from a memoryless learning tree for standard purity testing.
@@ E.2.3 Noise-robust detection using joint measurements
As in the lower bound, we utilize a basic reduction to the standard purity-testing problem to establishes our $O(1)$ two-copy upper bound. The first step is to use the intrinsic error-correction properties of the tensor network. In particular, [happy_2015] proposes a simple hierarchical majority-vote strategy for decoding the boundary state, known as the greedy decoder. While less robust than other decoding strategies, the greedy decoder allows us to obtain rigorous decoding guarantees. In particular, [happy_2015] gives us the following lemma.
@@
Under an erasure channel of strength $\lambda$, each out-leg of the boundary state is corrupted independently with probability $\lambda$. [happy_2015] shows that when the greedy decoder is applied to $t$ radial layers of the tensor network, the probability that an erasure error has propagated to a particular leg at radius RtR-t is $\leq\frac{1}{12}(12\lambda)^{\varphi^{R-t}}$. Applying a union bound over all legs at radius $r$, which we have shown is
@@
Hence, given two copies of $\mathcal{E}_{\lambda}(\rho_{b})$, we can apply the greedy decoder to both, and with probability $\geq 1-O(\exp\exp(-r(c-1)))$, we obtain two copies of $\rho$. In our problem, $\rho$ is either maximally mixed or a Haar-random pure state $\lvert\psi\rangle\!\langle\psi\rvert$. Our problem is then reduced to the well-studied purity testing problem, for which there is a simple two-copy algorithm. One can simply run a SWAP test $\textsf{SWAP}(\rho,\rho)$ (e.g., as described in [montanaro2018surveyquantumpropertytesting]); on two arbitrary states $\rho,\sigma$, the test accepts with probability